OpenStudy (sighn0more):

log(base)5 8 - log(base)5 (6-x) = log(base)5 x My textbook writes that the answer to this problem is x = 4 or 2. How would I go about getting this answer?

OpenStudy (freckles):

do you know the quotient rule for log thingys ?

OpenStudy (freckles):

\[\log_a(x)-\log_a(y)=\log_a(\frac{x}{y})\]

OpenStudy (sighn0more):

Yep! So far i know that it's log5 ( 8/(6-x) )

OpenStudy (freckles):

cool so you have: \[\log_5(\frac{8}{6-x})=\log_5(x) \implies \frac{8}{6-x}=x\]

OpenStudy (freckles):

now you could multiply both sides by (6-x)

OpenStudy (freckles):

8=x(6-x)

OpenStudy (freckles):

you have a quadratic equation to solve expand

OpenStudy (freckles):

and put into ax^2+bx+c=0 you can try to factor or use the quadratic formula or complete the square or whatever

OpenStudy (sighn0more):

Ohhh, I see now. I'll work the rest of this out. Thank you!

OpenStudy (freckles):

k